OpenStudy (anonymous):

Medal 24 / 3^2 / 8-5 - (-5)

OpenStudy (babynini):

Are those each individual questions?

OpenStudy (babynini):

or are those divided signs?

OpenStudy (anonymous):



OpenStudy (babynini):

Ok, let's start with the middle part :) 3^2 = ?

OpenStudy (anonymous):

9

OpenStudy (babynini):

8-5 = ?

OpenStudy (anonymous):

3

OpenStudy (babynini):

perfect! so now your equation looks like this\[24 \div \frac{ 9 }{ 3 }-(-5)\]

OpenStudy (babynini):

what is 9/3?

OpenStudy (anonymous):

3

OpenStudy (babynini):

fabulous! so now your equation is \[24\div3-(-5)\]

OpenStudy (babynini):

and if you simplify that its 24/3 + 5 Can you figure it out from here? :)

OpenStudy (anonymous):

13